OpenStudy (anonymous):

Compute (g o f )(x), given the functions below: f(x) = sqrt(3x-4) g(x) = 5x-8

OpenStudy (anonymous):

\[f(g(x))=f(5x-8)=\sqrt{3(5x-8)-4}\]

myininaya (myininaya):

satellite i forgot im suppose to be giving you medals sorry i been slacking

OpenStudy (anonymous):

of course you are supposed to simplify this \[f(g(x))=\sqrt{15x-24-4}=\sqrt{15x-28}\]

OpenStudy (anonymous):

one for you!

OpenStudy (anonymous):

g(f(x))=\[5\sqrt{3x-4}-8\]

myininaya (myininaya):

:)

OpenStudy (anonymous):

i don't get a medal because i computed \[f(g(x))\] when i t asked for \[g(f(x))\]

OpenStudy (anonymous):

brackett is right and i am wrong. btw is it samual brackett?

OpenStudy (anonymous):

can i get a medal ? ;) :D

OpenStudy (anonymous):

got it!

OpenStudy (anonymous):

nope i'm just brackett :)

OpenStudy (anonymous):

i got it right! so the tick was to spot the g of f!